OpenStudy (anonymous):

The length of a rectangle is twice its width. If the perimeter of the rectangle is 48ft , find its area.

OpenStudy (nurali):

Perimeter of rectangle = P = 48ft Length = 2w Width = w Perimeter of rectangle = 2(Length+Width) 48=2(2w+w) 48=2(3w) 48=6w 48/6=6w/6 8=w w=8ft Length = 2w = 2(8)=16ft Width = w = 8ft Area of rectangle = Length * Width Area of rectangle = 8 * 16 Area of rectangle = 128 square feet.